# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ 6.5 Adjoint System Specification
The fractional differential equations outlined below are satisfied by the adjoint variables:
@@
Proof.  The calculation of the partial derivative of the Hamiltonian yields the equations of the adjoint each corresponding to the state variable of consideration. For $\lambda_{S}$, we have:
@@
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
Thus, (28) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$. For $\lambda_{V}$, the derivative gives:
@@
$+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
which gives (29) under $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$. The remaining adjoint equations (30)–(35) can be derived in a similar manner by calculating the respective partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and using the fundamental relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$. The arrangement continues to embody the epidemiological relations where the transmission mechanisms are influenced by the cosate differentials and the passage rates determined by the compartments. $\square$
@@
Each part of the adjoint system can be seen recursively decomposed in the sequence structured by the equations $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$, which become independent once the other adjoint variables are fixed. This characteristic feature of the system corresponds to the compartmentalized configuration of the epidemiological dynamics and facilitates the computation of an efficient numerical solution.
6.6 Numerical Results of Optimal Control
Figure 7: Optimal Control Analysis : (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
Using Python, NumPy, and SciPy, the faster forward-backward sweep algorithm achieved convergence, exhibiting epidemic containment (peak: 170.0, final: 0.0) in 50 iterations within 30 seconds, and at a cost of 8750\$ Of the burial controls, which were the most cost-effective at 4\$/case, treatment with a 92% reduction in transmission, sequentially distilled intervention timing.
7 Efficiency Evaluation of Optimal Interventions
The numerical simulation of the fractional-order Ebola system employs the RK45 method. This method features adaptive step-size control which allows for a proper trade-off between precision and computational efficiency.
Algorithm 1 Runge-Kutta-Fehlberg (RKF45)
[…]
3:Initialize: $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
4:Set parameters: $h_{min}\leftarrow 10^{-8}$, $h_{max}\leftarrow 0.1$, $h\leftarrow 0.01$
@@ 7 Efficiency Evaluation of Optimal Interventions
6:  Compute RK coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$
7:  Calculate 4th/5th order solutions $\mathbf{y}_{4}$, $\mathbf{y}_{5}$
8:  Estimate error $\delta=\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
9:  Compute optimal step $h_{opt}=0.9h(\epsilon_{rel}/\delta)^{0.2}$
10:  if $\delta\leq\epsilon_{rel}$ then
11:   Accept step: $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
12:   Update $h\leftarrow\min(h_{max},\max(h_{min},h_{opt}))$
@@
14:   Reject step: $h\leftarrow\max(h_{min},h_{opt})$
@@
17:return Interpolated solution
@@ 7.1 Results and Discussion of Intervention Strategies
Figure 9: Foundational epidemiological principles regarding the impact of the treatment intervention
Figure 10: Effectiveness assessment of treatment strategy
@@
Figure 12: Comparison of key metrics for safe burial intervention.
Figure 13: Analysis of Safeguard Burial Approach Efficiency
@@
Figure 15: Key Metrics for the combined treatment and vaccination
Figure 16: Effectiveness assessment of combined vaccination and treatment
@@
Figure 18: Key metrics comparison for treatment-safe burial combination
Figure 19: Effectiveness assessment of treatment and safe burial strategy
@@
Figure 21: Key metrics for personal protection, vaccination, and treatment combination
Figure 22: Effectiveness assessment of three-intervention strategy
@@
Figure 24: Key metrics comparison for clinical and preventive combination
Figure 25: Effectiveness assessment of vaccination, treatment, and safe burial strategy
@@
Figure 27: Key metrics for comprehensive intervention approach
Figure 28: Effectiveness assessment of comprehensive strategy
7.2 Discussion
Evaluating strategies for controlling pandemics has unequally advanced in mortality reduction and has repeatedly failed in controlling transmission. In each case studied, no intervention profoundly altered the fundamental dynamics of spread, evidenced by the persistent values for peak infection rates and, in controlling the intervention to the baseline, a net nonexistence of infection post intervention. Mortality figures, on the other hand, were markedly disparate with strategies demonstrating either complementary or overlapping impacts when aggregated.
The recorded treatment intervention ($u_{3}\neq 0$) from integration of treatment with other controlling appliances yielded a 60.4% decrease in final mortality, and the safe burial ($u_{4}\neq 0$) showed 73.8% reduction in final mortality, thereby demonstrating the safe burial of deceased individuals treatment in burial both alone and in combination. The rationale for the incorporated benefits from Figures 1013 compartmental dynamics are on a shift in disease with no less than transmission chain breakage which possibly led to the two outcomes being flawed.
There were significant interactions when different types of interventions were used together. From figures $16$ vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive. In contrast, treatment with safe burial ($u_{3}\neq 0,u_{4}\neq 0$) and burial leads to 86.5% mortality reduction, which is a synergistic over additive effect that far exceeds the efficacy of the single interventions. The complementary effect are shown in the dynamics of Figure1919 as treatment drives the disease to fatal outcomes while safe burial prevents transmission after death.
@@ 7.2 Discussion
Combining interventions produced the same results. the combined three type intervention ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair. It is the combination of these results, documented in Figures 2225 and  $28$ that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective. In the context of EVD, personal protection does not mean social distancing; rather, it refers to the use of barrier techniques for those who come into direct interaction with infected persons or dead bodies. The combination of safe burial and treatment is synergistic and remains the optimal strategy for mortality reduction. It is the persistent inability to curtail infection prevalence that most strongly implies the mortality reduction is the consequence of the disease, postulated to contain altered outcome probabilities. This underscores the importance of post-exposure treatment combined with the clinical burial of the dead to reduce casualties.
8 Intelligent Disease Modeling – DINN of EVD
This segment presents the mathematical concepts supporting the architecture of the Disease-Informed Neural Network (DINN) as per the formulae of the fractional-order Ebola model that was discussed. (2). In the case of the DINN architecture, the fractional differential equations constrain the architectural configuration of the neural-embedded network, enabling the network to incorporate the differential equations of the epidemic system defined by the fractional-order differential equations, thus capturing the system’s physically plausible dynamics.
The governing system of fractional-order differential equations is expressed in the compact form:
$^{C}D_{t}^{\alpha}Y(t)+\mathcal{G}(Y(t;\chi))=0,\quad t\in[t_{0},T],$ (36)
The state Y(t) = [S(t), V(t), E(t), $I_{s}(t)$, $I_{a}(t)$, H(t), D(t), R(t)] along with the vector concept that every compartment population tends to zero. Let $\mathcal{G}(Y)=[g_{1}(Y),\ldots,g_{8}(Y)]$, where each function $g_{i}(Y)$ captures the heterogeneously nonlinear connections that inflow and outflow through its compartment. $\chi\in\mathbb{R}^{p}$ the parameters in the system which determines epidemiological corresponding with the degree/level of the interventions effectiveness and the intervention measures being undertaken. The neural network is approximated as centralized with the equations $NN_{w,b}(t):\mathbb{R}\longrightarrow\mathbb{R}^{8}$ which is initial feasible $Y(t)$ solution because of the abstract layer termed trainable beings the connected parameters the neural network in the $w$ and $b$ configurations as edges as nodes as the two fundamental elements in the neural network. The subsequent component of the function is termed the data portion is given by
$\mathcal{L}_{data}=\frac{1}{m}\sum_{s=1}^{m}\|NN_{w,b}(t_{s})-Y_{s}\|^{2}.$ (37)
The physical consistency is enforced through the residual term, which measures the discrepancy between the neural network output and the governing fractional dynamics:
$\mathcal{R}(NN_{w,b},t;\chi)={}^{C}D_{t}^{\alpha}NN_{w,b}+\mathcal{G}(NN_{w,b}(t)).$ (38)
The corresponding residual loss is computed as:
$MSE_{w,b,\chi}^{\mathcal{G}}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{R}(NN_{w,b},t_{i};\chi)\|^{2},$ (39)
The use of collocation points $t_{1},t_{2},\ldots,t_{N}$ discretizes the temporal domain. The entirety of this optimization problem simultaneously evaluates the parameters defining the neural network and the system by means of the composite objective:
@@ 8 Intelligent Disease Modeling – DINN of EVD
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(\mathcal{L}_{data}+\lambda MSE_{w,b,\chi}^{\mathcal{G}}\right),$ (40)
where $\lambda$ balances the contribution of data compliance and physical consistency. This formulation guarantees that the acquired solution serves the purpose of not only fitting the data but also complying with the core mechanistic precepts embedded in the fractional-order Ebola transmission model.
8.1 ML-EVD model
This subsection describes the concrete realization of the Disease-Informed Neural Network (DINN) framework design for the fractional-order Ebola epidemic model. The advanced physical consistency is maintained through an elaborate residual term that contains the entire system dynamics:
@@ 8.1 ML-EVD model
$\mathcal{R}(NN_{w,b},t;\chi)=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (41)
where the time-dependent force of infection is defined as:
$\lambda(t)=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$ (42)
Let $N$ denotes the total living population such that $N=S+V+E+I_{s}+I_{a}+H+R$. The residual loss function can be split into compartment-specific components, each measuring the gap between the neural network’s fractional derivatives and the corresponding dynamical equations:
$MSE_{residual}$ $=\sum_{X\in\mathcal{C}}MSE_{X},$ (43)
$MSE_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\right|^{2},$ (44)
@@
$MSE_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\right|^{2},$ (45)
@@
$MSE_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\right|^{2},$ (46)
@@
$MSE_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right|^{2},$ (47)
$MSE_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\right|^{2},$ (48)
$MSE_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\right|^{2},$ (49)
$MSE_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\right|^{2},$ (50)
$MSE_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\right|^{2},$ (51)
where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ indicates all the epidemiological compartments. Consistency with the observational measurements is guaranteed with the components set within the data adherence phase
$MSE_{data}$ $=\sum_{X\in\mathcal{C}}MSE_{X}^{data},$ (52)
$MSE_{X}^{data}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{o}|^{2},\quad X\in\mathcal{C},$ (53)
Let $X_{i}^{o}$ be the observed value of compartment $X$ at time $t_{i}$. The entire optimization task finds the values of both the neural network parameters and the parameters of the epidemiological model simultaneously:
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(MSE_{data}+MSE_{residual}\right),$ (54)
The collection of parameters $\chi=\{\Lambda,\mu,c(t),\beta,\eta_{a},\eta_{d},v(t),\varepsilon,\omega,\sigma,p,h_{s},\gamma_{s},\gamma_{a},\gamma_{h},\delta_{s},\delta_{h},\mu_{d},\alpha\}$ includes all constants related to the flow of epidemiological modeling to describe the transmission, progression of the disease and the impact of different control measures. Using this approach ensures that the learned solution integrates the observed data with the mechanistic underpinning of the Ebola transmission model.

### v3

@@ 5 Optimal Control Analysis
The adjoint variables satisfy the following system of fractional differential equations
@@
Proof.  The adjoint equations are derived by computing the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to each state variable, then applying the relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$ from Pontryagin’s maximum principle.
For $\lambda_{S}$, we compute
@@
$\quad+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
Equation (62) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$.
For $\lambda_{V}$, we have
@@
$\quad+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
which yields (63) via $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$.
The remaining adjoint equations (64)–(69) are obtained similarly by computing the corresponding partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and applying the fundamental adjoint relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$. Each derivative accounts for both direct effects of the state variable on the Hamiltonian and indirect effects through the force of infection $\lambda(t)$. $\square$
@@
The adjoint system presents a recursive structure, the equations for $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$ become independent once the other adjoint variables are determined. This structure reflects the compartmental organization of the epidemiological dynamics and facilitates efficient numerical solution of the optimality system.
5.1 Cost-Effectiveness of Optimal Control
Figure 4: Optimal Control Analysis (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
We performed numerical simulations in Python using the NumPy and SciPy libraries. After 50 iterations, the quick forward-backward sweep method achieved the ideal solution in 30 seconds. The optimal control strategy achieved epidemic containment, with the infected population peaking at 170 individuals before declining to zero, at a total intervention cost of 8750\$. Safe burial was the most cost-saving intervention, having a unit cost of 4\$ per case, and treatment resulted into a 92% proportion of reduction in transmission. The algorithm also identified the ideal time when each intervention should be put in place.
5.2 Numerical Assessment of intervention Strategies.
The RungeKutta45 scheme is used to perform the numerical analysis of the fractional-order Ebola system. This method uses the adaptive step-size control in order to balance the computational cost with the number of numerical accuracy needed to derive the best intervention strategy.
Algorithm 1 Runge-Kutta-Fehlberg (RKF45) Method for Adaptive Numerical Integration
1:Time interval $t_{0},T]$, initial state $\mathbf{y}_{0}$, system parameters $\mathbf{p}$, relative tolerance $\epsilon_{\text{rel}}=10^{-6}$, absolute tolerance $\epsilon_{\text{abs}}=10^{-8}$
2:High-precision numerical solution $\mathbf{y}(t)$ for $t\in[t_{0},T]$
3:Initialize $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
4:Set step size bounds $h_{\min}\leftarrow 10^{-8}$, $h_{\max}\leftarrow 0.1$, initial $h\leftarrow 0.01$
@@ 5.1 Cost-Effectiveness of Optimal Control
6:  Compute Runge-Kutta coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$ via standard RKF45 formulas
7:  Calculate 4th-order ($\mathbf{y}_{4}$) and 5th-order ($\mathbf{y}_{5}$) solutions
8:  Estimate local truncation error $\delta\leftarrow\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
9:  Compute optimal step size $h_{\text{opt}}\leftarrow 0.9h\left(\dfrac{\epsilon_{\text{rel}}\|\mathbf{y}_{5}\|+\epsilon_{\text{abs}}}{\delta}\right)^{0.2}$
10:  if $\delta\leq\epsilon_{\text{rel}}\|\mathbf{y}_{5}\|+\epsilon_{\text{abs}}$ then
11:   Accept step $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
12:   Update step $h\leftarrow\min(h_{\max},\max(h_{\min},h_{\text{opt}}))$
@@
14:   Reject step $h\leftarrow\max(h_{\min},h_{\text{opt}})$
@@
17:return Interpolated solution with 4th-order accuracy and 5th-order error control
@@ 5.3 Results and Discussion of Intervention Strategies
Evaluating strategies for controlling pandemics has unequally advanced in mortality reduction and has repeatedly failed in controlling transmission. In each case studied, no intervention profoundly altered the fundamental dynamics of spread, evidenced by the persistent values for peak infection rates and, in controlling the intervention to the baseline, a net nonexistence of infection post intervention. Mortality figures, on the other hand, were markedly disparate with strategies demonstrating either complementary or overlapping impacts when aggregated. The recorded treatment intervention ($u_{3}\neq 0$) from integration of treatment with other controlling appliances yielded a 60.4% decrease in final mortality, and the safe burial ($u_{4}\neq 0$) showed 73.8% reduction in final mortality, thereby demonstrating the safe burial of deceased individuals treatment in burial both alone and in combination. The rationale for the incorporated benefits from Figures 5 compartmental dynamics are on a shift in disease with no less than transmission chain breakage which possibly led to the two outcomes being flawed. There were significant interactions when different types of interventions were used together. From figures 7 vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive. In contrast, treatment with safe burial ($u_{3}\neq 0,u_{4}\neq 0$) and burial leads to 86.5% mortality reduction, which is a synergistic over additive effect that far exceeds the efficacy of the single interventions. The complementary effect are shown in the dynamics of figure 8 as treatment drives the disease to fatal outcomes while safe burial prevents transmission after death. Combining interventions produced the same results. The combined three type intervention in figure 9 ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair. It is the combination of these results, documented in figures 10 and  11 that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective. In the context of EVD, personal protection does not mean social distancing; rather, it refers to the use of barrier techniques for those who come into direct interaction with infected persons or dead bodies. The combination of safe burial and treatment is synergistic and remains the optimal strategy for mortality reduction. It is the persistent inability to curtail infection prevalence that most strongly implies the mortality reduction is the consequence of the disease, postulated to contain altered outcome probabilities. This underscores the importance of post-exposure treatment combined with the clinical burial of the dead to reduce casualties.
6 Disease-Informed Neural Network (DINN) for Ebola Modeling
This section introduces the Disease-Informed Neural Network (DINN), a physics-informed neural network framework that incorporates the fractional-order Ebola model (2) as physical constraints to ensure epidemiological consistency in predictions.
The fractional-order system governing Ebola dynamics is expressed as
$^{C}D_{t}^{\alpha}\mathbf{Y}(t)+\mathcal{G}(\mathbf{Y}(t);\bm{\chi})=0,\quad t\in[t_{0},T],$ (70)
where $\mathbf{Y}(t)=[S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t)]^{\top}$ denotes the state vector and $\bm{\chi}$ represents the epidemiological parameters.
A neural network $\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t):\mathbb{R}\to\mathbb{R}^{8}$ approximates the solution $\mathbf{Y}(t)$, where $\mathbf{w}$ and $\mathbf{b}$ are the network weights and biases. The training process minimizes two loss components: a data fidelity term and a physics consistency term
$\mathcal{L}_{\text{data}}$ $=\frac{1}{m}\sum_{s=1}^{m}\|\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{s})-\mathbf{Y}_{s}\|^{2},$
$\mathcal{L}_{\text{physics}}$ $=\frac{1}{N}\sum_{i=1}^{N}\|{}^{C}D_{t}^{\alpha}\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{i})+\mathcal{G}(\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{i});\bm{\chi})\|^{2}.$
Here, $\mathcal{L}_{\text{data}}$ penalizes deviations from observed data, while $\mathcal{L}_{\text{physics}}$ enforces the governing equations of the fractional-order model.
The DINN framework solves the composite optimization problem
@@ 6 Disease-Informed Neural Network (DINN) for Ebola Modeling
$(\mathbf{w}^{*},\mathbf{b}^{*},\bm{\chi}^{*})=\arg\min_{\mathbf{w},\mathbf{b},\bm{\chi}}\left(\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}\right),$ (71)
where $\lambda>0$ is used to balance between data fidelity and physical constraints. The suggested methodology will both optimize the weights of the neural network and estimate the epidemiological parameters, thus giving both a solution that is both in line with the empirical data and epidemiologically plausible.
6.1 DINN Architecture for the Ebola Model
In this subsection, the architecture of Disease-Informed Neural Network (DINN) that is used in the fractional-order Ebola model is described in detail. The network applies physical consistency directly with its incorporation of the model dynamics into the loss function.
The remaining vector that measures the difference between the outcome of the neural network and underlying epidemiological dynamics is denoted by the residual as
@@ 6.1 DINN Architecture for the Ebola Model
$\mathcal{R}(\mathcal{NN}_{\mathbf{w},\mathbf{b}},t;\bm{\chi})=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (72)
where the time-dependent force of infection is given by
$\lambda(t)=\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$
and the total living population is $N=S+V+E+I_{s}+I_{a}+H+R$.
The physics-informed loss function is decomposed into compartment-specific mean squared errors
$\mathcal{L}_{\text{residual}}$ $=\sum_{X\in\mathcal{C}}\mathcal{L}_{X},$
$\mathcal{L}_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\bigr|^{2},$
@@
$\mathcal{L}_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\bigr|^{2},$
@@
$\mathcal{L}_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\bigr|^{2},$
@@
$\mathcal{L}_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\bigr|^{2},$
$\mathcal{L}_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\bigr|^{2},$
$\mathcal{L}_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\bigr|^{2},$
$\mathcal{L}_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\bigr|^{2},$
$\mathcal{L}_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\bigr|^{2},$
where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ denotes the set of all epidemiological compartments.
Consistency with observed data is enforced by the data fidelity loss
$\mathcal{L}_{\text{data}}$ $=\sum_{X\in\mathcal{C}}\mathcal{L}_{X}^{\text{data}},$
$\mathcal{L}_{X}^{\text{data}}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{\text{obs}}|^{2},\quad X\in\mathcal{C},$
where $X_{i}^{\text{obs}}$ represents the observed value of compartment $X$ at time $t_{i}$.
The DINN framework solves the optimization problem
$(\mathbf{w}^{*},\mathbf{b}^{*},\bm{\chi}^{*})=\arg\min_{\mathbf{w},\mathbf{b},\bm{\chi}}\left(\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{residual}}\right),$ (73)
where the parameter vector $\bm{\chi}=\{\beta,\sigma,\gamma_{s},\eta_{a},\eta_{d},\varepsilon,v,\alpha\}$ contains key epidemiological constants governing transmission, progression, and control measures. This articulation guarantees that the solutions that are derived to fit both meet the observed data requirements and appear as per the mechanistic framework characteristic in the Ebola model.